The following figuTheres shows the profit of a green grocer for the past five years: R36 000,R550 000,R200 000,R80 000,and R700 000.The arithmetic mean of the data is:


1
Expert's answer
2021-07-30T14:13:47-0400

Arithmetic mean"=\frac {X_1+X_2+X_3+X_4+X_5} {n}"


Arithmetic Mean "=\frac {36,000+550,000+200,000+80,000+700,000}{5}"


Arithmetic Mean ="\frac {1,566,000}{5}"


Arithmetic Mean ="R313, 200"